# Defining implication relation for classical logic

### v11

@@ 3 Prove $(x\leftrightarrow y)\wedge x\leq y$;
    IRLA-BL0 relation. IRLA contains $x\leq y\Rightarrow x\to y=1$, i.e. $IRLA=IRLA+\{x\leq y\Rightarrow x\to y=1\}$; $\vdash_{BL0}IRLA$, so $BL0+\Delta=BL0$. Thus
@@
    IRLA-BL1 relation. IRLA contains $\{x\to y\leq\neg x\vee y\}$, i.e. $IRLA=IRLA+\{x\to y\leq\neg x\vee y\}$; $\vdash_{BL1}IRLA$, so $BL1+\Delta=BL1$. Thus
@@ 5.3 IRL-CL relationship
In $BL0(X,1,0,\wedge,\vee,\to,\leftrightarrow)$, if there exists $a\in X$ such that $a\neq 1$ and $a\neq 0$, then $\neg x\vee y\leq x\to y$ is not a theorem of BL0.
@@
From $a\neq 1$, $1\nleq a$, so $01\to a=$ 0 by definition; similarly, from $a\neq 0$, $a\to 0=0$. Let $t(x,y)=\neg x\vee y\to(x\to y)$. Then $t(1,a)=\neg 1\vee a\to(1\to a)=0\vee a\to 0=a\to 0=0\neq 1$, so $t(x,y)=1$ is not a theorem since its instance $t(1,a)\neq 1$. ∎
@@
In $IRLA(X,1,0,\wedge,\vee,\leftrightarrow,\to)$, $\neg x\vee y\leq(x\to y)$ is a theorem if and only if there exists no $a\in X$ such that $a\neq 1$ and $a\neq 0$.
@@
Necessity. If there exists $a\in X$ such that $a\neq 1$ and $a\neq 0$, then $t(x,y)=1$ is not a theorem of BL0 in terms of Lemma 5.3.1, neither is it a theorem of IRLA, since $\vdash_{BL0}IRLA$.
Sufficiency. If there exists no $a\in X$ such that $a\neq 1$ and $a\neq 0$, then for all $x,y\in X$, $x,y\in\{1,0\}$. In IRLA, $t(1,1)=\neg 1\vee 1\to(1\to 1)=1\to 1=1$, $t(1,0)=\neg 1\vee 0\to(1\to 0)=0\to 0=1$, $t(0,1)=\neg 0\vee 1\to(0\to 1)=1\to 1=1$, and $t(0,0)=\neg 0\vee 0\to(0\to 0)=1\to 1=1$. Thus, for all $x,y\in X$, $\neg x\vee y\to(x\to y)=1$. ∎
@@
Propositions are in the three types: tautology, contradiction and contingency.
@@
Both IRL and CL always contain tautologies (e.g. $\phi\vee\neg\phi$) and contradictions (e.g. $\phi\wedge\neg\phi$). So the only difference between the general case and the no-contingency is whether or not the system contains a contingency.
@@
Thus, for IRL and CL, it is the general case if and only if there exists a primitive contingency in the system, and it is the no-contingency case if and only if there exists no primitive contingency so that any formula is either tautological or contradictory.
For example, suppose the set of primitive propositional symbols (the signature of propositional language) is just the singleton $\{A\}$, and let $S_{A}$ be a system of IRL or CL with this signature. Then: (1) $S_{A}$ contains all the three types of propositions if neither $A$ nor $\neg A$ is an axiom; (2) $S_{A}$ contains no contingencies (each formula is either a tautology or a contradiction) if either $A$ or $\neg A$ is an axiom.
When not specifying a case in context, we mean it is in the general case.
@@
thus IRL is a sub-system of CL so that IRL is consistent (relative to CL), and the negation of Disjunction-to-Implication ¬($\neg(\neg\phi\vee\psi\to(\phi\to\psi))$ is not a theorem of IRL.
  • In the general case (i.e. there is a primitive contingency):
@@
  • In the no-contingency case (i.e. there is no primitive contingency):
    $IRL=CL\textup{, and }\neg\phi\vee\psi\to(\phi\to\psi)\textup{ is a theorem of }IRL.$
@@
In the no-contingency case, any formula is either tautological or contradictory. However, it does not mean that the system is syntactically complete in this case, because we only know that every primitive proposition is a tautology or contradiction, but we do not know whether it is a tautology or a contradiction, and this is underivable.
@@ 5.4 Provable-equivalence class of Disjunction-to-Implication
    $\phi\to(\top\to\phi)$, $\neg\phi\to(\phi\to\bot)$. (Minimum of univariate implication)
@@
Note: Each of (1) and (2) is a pair of dual formulas that are provably equivalent in IRL.
@@
(1) Minimum of univariate implication $\to$ (2) Generalized ECQ:
1 $\phi$ Premise 2 $\top\to\phi$ 1: Minimum of univariate implication 3 ϕitalic-$\phi\to\top$ Dual of ECQ 4 ϕitalic-$\phi\leftrightarrow\top$ 2,3: Bi-implication introduction 6 $\psi\to\phi$ 4: Safe generalized ECQ
@@
(6) Assertion $\to$ (1) Minimum of univariate implication:
@@
(9) Top-or-bottom $\to$ (1) Minimum of univariate implication:
1 $(\phi\to\bot)\to\neg\phi$ Maximum of univariate implication 2 $(\phi\to\bot)\vee(\top\to\phi)\to\neg\phi\vee(\top\to\phi)$ 1: Compatibility of $\to$ with $\vee$ 3 $(\phi\to\bot)\vee(\top\to\phi)\leftrightarrow\top$ Top-or-bottom 4 $\top\to\neg\phi\vee(\top\to\phi)$ 3,2: Replacement 5 $\phi\to(\top\to\phi)$ 4: Definition of implication
@@
1 $(\top\to\phi)\vee(\top\to\neg\phi)$ Conditional excluded middle 2 $(\top\to\phi)\vee(\phi\to\bot)$ Contraposition,1
@@
Chain of proofs:
Generalized ECQ (known) $\to$ Conjunction-to-equivalence (new) $\to$ Minimum of univariate implication (known).
@@
Conjunction-to-equivalence $\to$ Minimum of univariate implication:
@@
6 Discussion
6.1 The decision problem
In this work, neither a decision procedure nor an undecidability proof for IRL is found.
A conjecture: IRL is undecidable. If IRL is indeed undecidable, it is a reflection of the nature of implication from the original meaning of implication relation.
Examination of (in)validity can be done on a case-by-case basis for a specific of formula (schema). The following facts and ideas are useful for the case-by-case approach.
@@ 6.1 The decision problem
    If a formula is valid in IRL, it must be valid in both classical logic CL and the contra-classical logic CL0 (cf. Subsection 5.3 and Remark 5.2.1); hence if a formula is invalid in either CL or CL0, it is invalid in IRL.
@@
    The (in)validity of a formula in CL can be systematically determined by a classical decision procedure such as truth table and truth tree.
@@
    Since $IRL\cong IRLA$, $CL0\cong BL0$, $CL\cong BL1$, both syntactic and semantic derivations can be done. For example, examining the invalidity of a schema $\eta(\phi,\psi)$ is equivalent to examining if $t(x,y)\neq 1$ for the corresponding term.
@@
    A schema is invalid if an instance of it is invalid. Let $\eta(\phi_{1},...,\phi_{n})$ be a schema and $t(x_{1},...,x_{n})$ its corresponding term. If there exists $(u_{1},...,u_{n})$ such that $t(u_{1},...,u_{n})\neq 1$ in IRLA, then $\eta$ is invalid in IRL; if there exists $(u_{1},...,u_{n})$ such that $t(u_{1},...,u_{n})\neq 0$ in IRLA, then $\neg\eta$ is invalid in IRL.
Example 6.1.1 (Invalidity of Peirce’s law).
Peirce’s law $((\phi\to\psi)\to\phi)\to\phi$ is invalid in IRL in the general case.
Let $t(x,y)=((x\to y)\to x)\to x$.
In the general case, there must be a contingency $a$, so $a\neq 1$ and $a\neq 0$.
In BL0, since $a\neq 1$ and $a\neq 0$, $01\to a=$ 0 and $a\to 0=0$. Hence
$t(a,0)$ $=((a\to 0)\to a)\to a$
$=(0\to a)\to a$
$=1\to a$
$=0.$
So, $t(x,y)=1$ is not a theorem of BL0 since its instance $t(a,0)=0\neq 1$. Thus, $t(x,y)=1$ is not a theorem of IRLA either, as IRLA is a sub-system of BL0.
Therefore, Peirce’s law is invalid in IRL.
On the other hand, in IRLA, $t(1,1)=((1\to 1)\to 1)\to 1=(1\to 1)\to 1=1\to 1=1$, indicating that $t(x,y)=0$ is not a theorem of IRLA since its instance $t(1,1)=1\neq 0$. In fact, as Peirce’s law is a theorem of CL, its negation can not be a theorem of IRL.
@@
Thus, Peirce’s law, as Disjunction-to-Implication, is independent of IRL. It can be shown that Peirce’s law is a provable consequence of Disjunction-to-Implication, but (very likely) not vice versa.
6.2 Interpretation of Generalized ECQ of implication
Generalized ECQ $\psi\to(\phi\to\psi)$ means that, for any propositions $\phi$ and $\psi$, if $\psi$ is true, then there exists a “certain mechanism” of implication from $\phi$ to $\psi$ and $\phi\to\psi$ is true, conflicting with the original meaning of implication by which implication is non-truth functional.
6.3 Interpretation of ECQ
A simple conditional proof of ECQ $\phi\wedge\neg\phi\to\psi$ is
1 $\phi$ Premise 2 $\neg\phi$ Premise 3 $\phi\vee\psi$ 1: Disjunction introduction 4 $\psi$ 3,2: Disjunctive syllogism
Hence ECQ is necessary for disjunction introduction and disjunctive syllogism – both do not conflict to the original meaning of implication. Therefore, ECQ contributes to the goal of strength for the system (cf. Subsection 2.6).
Furthermore, ECQ $\bot\to\varphi$ and its dual $\varphi\to\top$ can be interpreted to intuition as “if an always-false thing is (was) true, then anything is (would be) true” and “in any cases, an always-true thing is true”, respectively, as similarly mentioned, e.g., in (Ceniza, 1988).
A “standard (set of rules for) interpretation” of ECQ may be explored for practical purpose. For example, given a conditional $P\to Q$: if $P$ is a contradiction (e.g. it is well known that $P$ is false), then $P\to Q$ can be interpreted as “if the false thing like $P$ were true, anything including $Q$ must be true”; if $Q$ is a tautology (e.g. it is a well known fact), then $P\to Q$ can be interpreted as “in any cases including $P$, an always true thing like $Q$ is true”.

### v12

@@ 3 Prove $(x\leftrightarrow y)\wedge x\leq y$;
Remark 5.2.2 (CL is a super-system of IRL).
Actually, the idea in the proof of the above Lemma 5.2.3 can be used in proving directly that classical propositional logic CL is a supper-system of IRL. This means that the consistency of IRL can be proved before any derivations.
@@
    IRLA-BL0 relation. $\vdash_{IRLA}\{x\leq y\Rightarrow x\to y=1\}$, i.e. $IRLA=IRLA+\{x\leq y\Rightarrow x\to y=1\}$; $\vdash_{BL0}\Delta$, i.e. $BL0+\Delta=BL0$. Thus
@@
    IRLA-BL1 relation. $\vdash_{IRLA}\{x\to y\leq\neg x\vee y\}$, i.e. $IRLA=IRLA+\{x\to y\leq\neg x\vee y\}$; $\vdash_{BL1}\Delta$, i.e. $BL1+\Delta=BL1$. Thus
@@ 5.3 IRL-CL relationship
In $BL0=(X,1,0,\wedge,\vee,\to,\leftrightarrow,\leq)$, if there exists $c\in X$ such that $c\neq 1$ and $c\neq 0$, then $\neg x\vee y\leq x\to y$ is not a theorem of BL0.
@@
From $c\neq 1$, $1\nleq c$, so $01\to c=$ 0 by definition; similarly, from $c\neq 0$, $c\to 0=0$. Let $t(x,y)=\neg x\vee y\to(x\to y)$. Then $t(1,c)=\neg 1\vee c\to(1\to c)=0\vee c\to 0=c\to 0=0\neq 1$, so $t(x,y)=1$ is not a theorem of BL0 since its instance $t(1,c)\neq 1$. ∎
@@
In $IRLA=(X,1,0,\wedge,\vee,\leftrightarrow,\to,\leq)$, $\neg x\vee y\to(x\to y)=1$ is a theorem if and only if there exists no $c\in X$ such that $c\neq 1$ and $c\neq 0$.
@@
Necessity. If there exists $c\in X$ such that $c\neq 1$ and $c\neq 0$, then $t(x,y)=1$ is not a theorem of BL0 in terms of Lemma 5.3.1, neither is it a theorem of IRLA, since $\vdash_{BL0}IRLA$.
Sufficiency. If there exists no $c\in X$ such that $c\neq 1$ and $c\neq 0$, then for all $x,y\in X$, either $x=1$ or $x=0$ (i.e. $x,y\in\{1,0\}$). In IRLA, $t(1,1)=\neg 1\vee 1\to(1\to 1)=1\to 1=1$, $t(1,0)=\neg 1\vee 0\to(1\to 0)=0\to 0=1$, $t(0,1)=\neg 0\vee 1\to(0\to 1)=1\to 1=1$, and $t(0,0)=\neg 0\vee 0\to(0\to 0)=1\to 1=1$. Thus, for all $x,y\in X$, $\neg x\vee y\to(x\to y)=1$, namely it is a theorem of IRLA in this case. ∎
@@
Propositions are in the three types: tautology, contradiction, and contingency.
@@
Both IRL and CL always contain tautologies (e.g. $\phi\vee\neg\phi$) and contradictions (e.g. $\phi\wedge\neg\phi$). So the only difference between the general case and the no-contingency case is whether or not the system contains a contingency.
@@
Thus, for IRL and CL, it is the general case if and only if there exists a primitive contingency in the system, and it is the no-contingency case if and only if there exists no primitive contingency i.e. any formula is either tautological or contradictory.
For example, suppose the set of primitive propositional symbols (the signature of propositional language) is simply the singleton $\{A\}$, and let $S_{A}$ be an instance of IRL or CL with this signature. Then: (a) $S_{A}$ contains all the three types of propositions if neither $A$ nor $\neg A$ is an axiom; (b) $S_{A}$ contains no contingencies (each formula is either a tautology or a contradiction) if either $A$ or $\neg A$ is an axiom.
When not specified, we mean it is in the general case.
@@
thus IRL is a sub-system of CL, so IRL is consistent (relative to CL), and the negation of Disjunction-to-Implication ¬($\neg(\neg\phi\vee\psi\to(\phi\to\psi))$ is not a theorem of IRL.
  • In the general case (the system contains all types of propositions including contingencies, which is equivalent to that there is a primitive contingency):
@@
  • In the no-contingency case (the system contains no contingencies, which is equivalent to that there is no primitive contingency):
    $IRL=CL\textup{, and }\neg\phi\vee\psi\to(\phi\to\psi)\textup{ is a theorem of the system.}$.
@@
In the no-contingency case, any formula is either tautological or contradictory. However, it does not mean that the system is syntactically complete in this case, because we only know that every primitive proposition is either a tautology or a contradiction, but we do not know which one it is, and this is underivable.
@@ 5.4 Provable-equivalence class of Disjunction-to-Implication
    $\phi\to(\top\to\phi)$, $\neg\phi\to(\phi\to\bot)$. (Lower bound of univariate implication)
@@
Note that each of 1 and 2 is a pair of dual formulas that are provably equivalent in IRL.
@@
(1) Lower bound of univariate implication $\to$ (2) Generalized ECQ:
1 $\phi$ Premise 2 $\top\to\phi$ 1: Lower bound of univariate implication 3 $\psi\to\phi$ 2: Safe generalized ECQ
@@
(6) Assertion $\to$ (1) Lower bound of univariate implication:
@@
(9) Top-or-bottom $\to$ (1) Lower bound of univariate implication:
1 $\phi$ Premise 2 $(\phi\to\bot)\to\neg\phi$ Upper bound of univariate implication 3 ¬(ϕ)italic-$\neg(\phi\to\bot)$ 1,2: Modus tollens 4 $(\top\to\phi)\vee(\phi\to\bot)$ Top-or-bottom 5 $\top\to\phi$ 3,4: Disjunctive syllogism
@@
1 ($(\top\to\neg\phi)\leftrightarrow(\phi\to\bot)$ Contraposition, Double negation 2 $(\top\to\phi)\vee(\top\to\neg\phi)$ Conditional excluded middle 3 $(\top\to\phi)\vee(\phi\to\bot)$ 1,2: Replacement
@@
It is done in the chain Generalized ECQ (known) $\to$ Conjunction-to-equivalence (new) $\to$ Lower bound of univariate implication (known):
@@
Conjunction-to-equivalence $\to$ Lower bound of univariate implication:
@@
6 The decision problem
Since $IRL\cong IRLA$ (cf. Subsection 4.5), a formula schema $F(\phi_{1},...,\phi_{n})$ in IRL corresponds to the term $t(x_{1},...,x_{n})$ in IRLA (cf. Table 4).
In a sentence $t(x_{1},...,x_{n})=1$ of an algebraic system with an underlying set $X$, all variables are implicitly universally quantified over $X$. So the decision problem considered in $IRLA=(X,1,0,\neg,\wedge,\vee,\leftrightarrow,\to,\leq)$ is, given any term $t(x_{1},...,x_{n})$, decide whether or not $t(x_{1},...,x_{n})=1$ for all $(x_{1},...,x_{n})\in X^{n}$ – we may simply say whether or not “$t(x_{1},...,x_{n})$ is valid” for convenience.
A complete decision procedure is not found for IRL/IRLA in this work. A conjecture: IRL/IRLA is undecidable. If it is indeed undecidable, this is a reflection of the nature of implication relation, which comes from the original meaning of implication.
If $t(x_{1},...,x_{n})$ is valid in IRLA, it is valid in both its super-systems BL0 and BL1 (cf. Subsection 5.2). So if $t(x_{1},...,x_{n})$ is invalid in either BL0 or BL1, it is invalid in IRLA. Thus, BL0 and BL1 can be used as two basic validity “filters” for IRLA. For example, ¬($\neg(\neg x\vee y\to(x\to y))$ is invalid in BL1, so it is invalid in IRLA.
Since BL1 represents classical propositional logic (cf. Subsection 5.2), many ready decision procedures can be utilized, such as the basic methods of truth table, truth tree, and SAT algorithms.
The following is a simple validity-checking method for BL0.
Validity checking in BL0 on a minimal set
Consider the validity-checking problem in $BL0=(X,1,0,\neg,\wedge,\vee,\to,\leq)$: given any expression $t(x_{1},...,x_{n})$, whether or not $t(x_{1},...,x_{n})$ is valid, i.e. if $t(x_{1},...,x_{n})=1$ for all $(x_{1},...,x_{n})\in X^{n}$?
Suppose there exists $c\in X$ such that $10<c<$ 1 (i.e. there exists a contingency in the system). Consider validity checking on the set $\{0,c,1\}$, a minimal subset of $X$ that can reflect the general case (there are all the three types of propositions: tautology, contradiction and contingency).
[…]
Table 6: Closure of the set $\{0,c,1\}$ in BL0 *
      $x$       $y$       $\neg x$       $x\wedge y$       $x\vee y$       $x\to y$
      00       00       $1$       00       00       $1$
      00       $c$       $1$       00       $c$       $1$
      00       $1$       $1$       00       $1$       $1$
      $c$       00       $\neg c$       00       $c$       00
      $c$       $c$       $\neg c$       $c$       $c$       $1$
      $c$       $1$       $\neg c$       $c$       $1$       $1$
      $1$       00       00       00       $1$       00
      $1$       $c$       00       $c$       $1$       00
      $1$       $1$       00       $1$       $1$       $1$
[…]
Thus, in BL0, any expression $t(x_{1},...,x_{n})\in\{0,c,\neg c,1\}$ for all $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$. Note that $\neg c$ is also a contingency, i.e. $10<\neg c<$ 1, since, only in a Boolean algebra, $10<x<$ 1 if and only if $10<\neg x<$ 1 for all $x$ in its underlying set.
If $t(x_{1},...,x_{n})\neq 1$ for some $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$, then $t(x_{1},...,x_{n})\neq 1$ for some $(x_{1},...,x_{n})\in X^{n}$, so the answer to the validity question of $t(x_{1},...,x_{n})$ is “No”.
Otherwise $t(x_{1},...,x_{n})=1$ for all $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$, the answer to the validity question of $t(x_{1},...,x_{n})$ is “Unknown”, i.e., it is possible for $t$ is valid and it is also possible for $t$ is invalid, since, although $t=1$ for all $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$, it is possible that $t\neq 1$ for some $(x_{1},...,x_{n})\in X^{n}\setminus\{0,c,1\}^{n}$.
In terms of these facts, we have a simple algorithm for the (in)validity-checking problem as described below.
Validity checking by exhaustion over $\{0,c,1\}\subset X$
In BL0:
@@ 00       $1$       $1$       00       $1$       $1$
    Input an expression $t(x_{1},...,x_{n})$;
@@
    For each $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$, compute $t(x_{1},...,x_{n})$;
@@
    At any step:
    • If $t(u_{1},...,u_{n})\neq 1$ (i.e. $t=0,c,\neg c$) for some $(u_{1},...,u_{n})\in\{0,c,1\}^{n}$, $t$ is Invalid with a countermodel $(x_{1},...,x_{n})=(u_{1},...,u_{n})$;
    • If $t(x_{1},...,x_{n})\neq 0$ (i.e. $\neg t\neq 1$) for some $(u_{1},...,u_{n})\in\{0,c,1\}^{n}$, $\neg t$ is invalid with a countermodel $(x_{1},...,x_{n})=(u_{1},...,u_{n})$;
    • If both $t$ and $\neg t$ are invalid, $t$ is independent of the system;
@@
    If $t(x_{1},...,x_{n})=1$ for all $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$, the validity of $t$ is Unkown (but $\neg t$ is invalid);
  5. 5.
    If $t(x_{1},...,x_{n})=0$ for all $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$, $t$ is Invalid (but the validity of $\neg t$ is unknown).
For IRLA as a sub-system of BL0:
  • $t$ is invalid in IRLA if it is invalid in BL0.
  • $t$ is independent of IRLA if it is independent of BL0.
  • The validity of $t$ is unknown in IRLA when its validity is unknown in BL0, or $t$ is valid in BL0.
This simple exhaustive search algorithm, with the worst-case complexity $3^{n}$, is of course incomplete, since it is not for all inputs that the algorithm can give “Yes” or “No” answer (for some inputs the answer is “Unknown”). But it is useful as a basic validity “filter” for IRLA/IRL.
Example 6.0.1.
Is $t(x)=(1\to x)\to x$ valid? For $x\in\{0,c,1\}$:
  • 1.
    $x=0$: $t(0)=(1\to 0)\to 0=0\to 0=1$.
  • 2.
    $x=c$: $t(c)=(1\to c)\to c=0\to c=1$.
  • 3.
    $x=1$: $t(1)=(1\to 1)\to 1=1\to 1=1$.
The result is $t(x)=(1\to x)\to x=1$ for all $x$ in $\{0,c,1\}$, so the answer to the validity question of $(1\to x)\to x$ is “Unknown” (i.e. $(1\to x)\to x$ is possibly valid and possibly invalid as well).
Example 6.0.2.
Is $t(x)=x\to(1\to x)$ valid? For $x\in\{0,c,1\}$:
  • 1.
    $x=0$: $t(0)=0\to(1\to 0)=0\to 0=1$.
  • 2.
    $x=c$: $t(c)=c\to(1\to c)=c\to 0=0$. (a countermodel)
So $t(x)=x\to(1\to x)$ is invalid with a countermodel $x=c$. And since $\neg t(0)=\neg 1=0$, $\neg t(x)$ is also invalid, thus $t(x)$ is independent of IRLA.
Example 6.0.3.
Is $t(x,y)=(x\to y)\to\neg x\vee y$ valid? For $(x,y)\in\{0,c,1\}^{2}$:
  • 1.
    $(x,y)=(0,0)$: $t(0,0)=(0\to 0)\to\neg 0\vee 0=1\to 1=1$.
  • 2.
    $(x,y)=(0,c)$: $t(0,c)=(0\to c)\to\neg 0\vee c=1\to 1=1$.
  • 3.
    $(x,y)=(0,1)$: $t(0,1)=(0\to 1)\to\neg 0\vee 1=1\to 1=1$.
  • 4.
    $(x,y)=(c,0)$: $t(c,0)=(c\to 0)\to\neg c\vee 0=0\to\neg c=1$.
  • 5.
    $(x,y)=(c,c)$: $t(c,c)=(c\to c)\to\neg c\vee c=1\to 1=1$.
  • 6.
    $(x,y)=(c,1)$: $t(c,1)=(c\to 1)\to\neg c\vee 1=1\to 1=1$.
  • 7.
    $(x,y)=(1,0)$: $t(1,0)=(1\to 0)\to\neg 1\vee 0=0\to 0=1$.
  • 8.
    $(x,y)=(1,c)$: $t(1,c)=(1\to c)\to\neg 1\vee c=0\to c=1$.
  • 9.
    $(x,y)=(1,1)$: $t(1,1)=(1\to 1)\to\neg 1\vee 1=1\to 1=1$.
The result is $t(x,y)=(x\to y)\to\neg x\vee y=1$ for all $(x,y)$ in $\{0,c,1\}^{2}$, so the answer to the validity question of $(x\to y)\to\neg x\vee y$ is “Unknown” (i.e. $(x\to y)\to\neg x\vee y$ is possibly valid and possibly invalid as well).
Example 6.0.4.
Is $t(x,y)=\neg x\vee y\to(x\to y)$ valid? For $(x,y)\in\{0,c,1\}^{2}$:
  • 1.
    $(x,y)=(0,0)$: $t(0,0)=\neg 0\vee 0\to(0\to 0)=1\to 1=1$.
  • 2.
    $(x,y)=(0,c)$: $t(0,c)=\neg 0\vee c\to(0\to c)=1\to 1=1$.
  • 3.
    $(x,y)=(0,1)$: $t(0,1)=\neg 0\vee 1\to(0\to 1)=1\to 1=1$.
  • 4.
    $(x,y)=(c,0)$: $t(c,0)=\neg c\vee 0\to(c\to 0)=\neg c\to 0=1\to c=0$. (a countermodel)
So $t(x,y)=\neg x\vee y\to(x\to y)$ is invalid with a countermodel $(x,y)=(c,0)$. And as we also have e.g. $t(0,0)=1$, $t(x,y)$ is independent of IRLA.
Example 6.0.5.
Are $t(x,y)=((x\to y)\to x)\to x$ and its negation $\neg t$ valid? For $(x,y)\in\{0,c,1\}^{2}$:
  • 1.
    $(x,y)=(0,0)$: $t(0,0)=((0\to 0)\to 0)\to 0=(1\to 0)\to 0=0\to 0=1$. (a countermodel for $\neg t$)
  • 2.
    $(x,y)=(0,c)$: $t(0,c)=((0\to c)\to 0)\to 0=(1\to 0)\to 0=0\to 0=1$.
  • 3.
    $(x,y)=(0,1)$: $t(0,1)=((0\to 1)\to 0)\to 0=(1\to 0)\to 0=0\to 0=1$.
  • 4.
    $(x,y)=(c,0)$: $t(c,0)=((c\to 0)\to c)\to c=(0\to c)\to c=1\to c=0$. (a countermodel for $t$)
So, $t(x,y)=((x\to y)\to x)\to x$ is invalid with a countermodel $(x,y)=(c,0)$, and $\neg t(x,y)$ is also invalid with a countermodel $(x,y)=(0,0)$. Thus $((x\to y)\to x)\to x$ is independent of IRLA.
@@
Example 6.0.5 is actually about the validity of Peirce’s law, and the result indicates that Peirce’s law, like Disjunction-to-Implication, is independent of IRL. It can be shown that, in IRL, Peirce’s law is a provable consequence of Disjunction-to-Implication, but (very likely) not vice versa.